[figure: 9.7 cm]

STUDY OF THE MOTT TRANSITION IN THE 3D HUBBARD

MODEL

F. Mancini, V. Turkowski Dipartimento di Scienze Fisiche ”E.R.Caianiello” e Unita I.N.F.M. di Salerno Università di Salerno, 84081 Baronissi (SA), Italy, IIASS ”E.R.Caianiello”, 84019 Vietri sul Mare (SA), Italy

The Mott-Hubbard transition in the 3D Hubbard model is studied using the composite operator method. The critical interaction U_{c}, which separates paramagnetic insulating and metallic phases, has been found. In order to investigate the ground state the probability of intersite hopping is studied. It is shown that some correlation functions go to zero as the interaction reaches the critical value U_{c}. This fact gives reason to consider these functions as the order characteristics in the normal phase.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : V. Turkowski

Address for correspondence : Dipartimento di Scienze Fisiche ”E.R.Caianiello” e Unita I.N.F.M. di Salerno Universita di Salerno, 84081 Baronissi (SA), Italy

Email address : turkowsky@vaxsa.csied.unisa.it